OpenStudy (anonymous):

Can someone help? Kelly went to a hockey game with her family. They bought 4 bags of cotton candy and 2 souvenir cups and spent $36. Marissa's family went to the same game, bought 7 bags of cotton candy and 3 souvenir cups, and spent $59. How much did one souvenir cup cost at the game? a. $6 b.$8 c.$5 d.$7

OpenStudy (anonymous):

3( 4c + 2s = 36 ) 2 ( 7c + 3s = 59 ) 12 c + 6 s = 102 14 c + 6 s = 118 -2c = - 16 c = 8